The parallel plates of an air capacitor are seperated by 2.25 mm. Each plate carries a charge of 6.5 nC. The magnitude of the electric field of the plates is 4.75x10⁵ V/m. Find the a) potential difference between the plates, b) capacitance, and c) area of a plate.
1
Expert's answer
2019-03-05T12:26:58-0500


a) By definition of the potential difference is work on moving point charge between two plates. So


"\Delta\phi=E\cdot d= 4.75\cdot10⁵ V/m \cdot 0.00225 m =1068.75 V"


b)capacitance


"C= \frac{Q}{\Delta\phi}=\frac{6.5 nC}{1068.75V}=6.08187\cdot10^{-12} F"


c )

"S=\frac{C\cdot d}{\varepsilon}=\frac{6.08187\cdot10^{-12} F\cdot 0.00225 m}{8.854F/m}=0.0015 m^2"

Answer

"\Delta\phi=1068.75 V""C=6.08187\cdot10^{-12} F"

"S=0.0015 m^2"